### v1

@@ Abstract
We present several equivalent conditions of the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ under mild assumptions, where $C(X)$ denotes the lattice of closed subsets of a $T_{0}$ topological space. We also provide an example of a non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous. Additionally, we show that a $T_{0}$ space is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets is a quasicontinuous (quasialgebraic) domain by using $n$-approximation. Furthermore, we provide a necessary condition for when a topological space possesses a Scott completion. This allows us to give more examples which do not have Scott completions.
@@ 1 Introduction
Topological semilattice has a deep connection with the theory of continuous lattices in domain theory. For instance, Lawson showed that compact unital semilattices with small semilattices are exactly continuous lattices with Lawson topology, see [11, Theorem VI-3.4]. From the viewpoint of category topology, Banaschewski introduced the notion of an essential embedding in the category of $T_{0}$ spaces and continuous maps [2], which has a close relation to Scott’s injective spaces. He also obtained a characterization of so-called essentially complete $T_{0}$ spaces in terms of join filters. Later Hoffmann provided a more transparent characterization: essentially complete $T_{0}$ spaces are sober spaces which are complete lattices in their specialisation order and satisfy an additional condition connecting the sup-operator with the topology [14]. In this paper we study a special kind of essentially complete $T_{0}$ space, namely the Scott space of the lattice of closed subsets of a topological space with the supremum operator.
Directed spaces [24] are special $T_{0}$ topological spaces, extending the concept of Scott spaces. The idea is that the priori is a $T_{0}$ space and its convergent directed subsets relative to the specialization order rather than a poset and its existing directed supremums. In $T_{0}$ topological spaces, the notion of directed spaces is equivalent to that of monotone determined spaces defined by Erné [8], also is equivalent to crp-spaces introduced by Battenfeld [3, Definition 6.7], in this paper we always assume that topological space is $T_{0}$. The category of directed spaces was shown to be Cartesian closed [24, 20]. $T_{0}$ c-spaces (or $n$-continuous spaces) are all monotone determined spaces. Based on this, Feng and Kou [9] introduced another approximation relation on monotone determined spaces, called $d$-approximation, and defined the quasicontinuous monotone determined spaces. They established an equivalence between quasicontinuous spaces and locally hypercompact spaces. Recently, Zhang, Shi and Li [25] have also contributed to this field by introducing the notions of continuous spaces and quasicontinuous spaces, focusing on these spaces within weak monotone convergence spaces.
@@
It is well known that the hull-kernel topology of the spectrum ${\mathrm{Spec}L}$ equals to the Scott topology when $L$ is a completely distributive (or distributive hypercontinuous) lattice, see [11]. There is a long-standing open problem asked by Lawson and Mislove that for which distributive continuous lattice its spectrum with the hull-kernel topology is exactly a sober locally compact Scott space (see [23, Problem 528]). Chen, Kou and Lyu [4] found a necessary condition for this problem, which is $\sigma(L^{op})=v(L^{op})$. Since $L^{op}$ can be viewed as a lattice of closed subsets of a topological space, we continue to investigate the equivalent conditions of $\sigma(C(X))=v(C(X))$ for general topological spaces $X$. Surprisingly the continuity of the supremum function $\Sigma C(X)\times\Sigma C(X)\rightarrow\Sigma C(X)$ is one of these equivalent conditions under mild assumptions. For monotone determined spaces, we can drop the extra assumption. In addition, we provide an example of non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous.
In domain theory, there is a close correspondence between the continuity of a dcpo $L$ and the continuity of the lattice of its Scott open subsets $\sigma(L)$. More surprisingly, the continuity of a dcpo $L$ can be characterized by the continuity of its lattice of its Scott closed subsets $\Gamma(L)$ (for example, see [4, Theorem 3.17]). These correspondences can be extended to general $T_{0}$ spaces. We show that a $T_{0}$ space $X$ is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets $C(X)$ is quasicontinuous (quasialgebraic).
@@
Monotone convergence spaces are very important in domain theory. Every $T_{0}$ space has a free monotone convergence completion called D-completion (for example, see [16]). We show that D-completion of any monotone determined space is a Scott space. The concept of Scott completion is introduced by Zhang, Shi and Li in [25], which has a slight difference with D-completion. We also obtain that Scott-completion of any monotone determined space coincides with its D-completion as a corollary. The Scott-completion of monotone determined spaces are shown to be very useful to construct free dcpo algebras over any general dcpo (see [6, 7]). But the notion does not seem to have been studied much. We provide a necessary condition for when a space has Scott completion. Based on this we give a simple proof of a previous result in [17, Theorem 4.1].
@@
Outline. In Section 2, we recall some useful knowledge about monotone determined spaces. We devote Section 3 to a discussion of D-completion of monotone determined spaces. In Section 4 the Scott space of the lattice of closed subsets of a topological space with the jointly continuous supremum operator is discussed. We characterize the quasicontinuity of lattice of closed subsets in Section 5. Finally more examples of topological spaces which do not have Scott completions are presented in Section 6.
@@ 2 Preliminaries
A nonempty set $P$ endowed with a partial order $\leq$ is called a poset. For $A\subseteq P$, we set $\downarrow A=\{x\in P:\exists a\in A,\ x\leq a\}$$\uparrow A=\{x\in P:\exists a\in P,\ a\leq x\}$$A$ is called a lower or an upper set, if $A=\downarrow A$ or $A=\uparrow A$ respectively. For an element $a\in P$, we use $\downarrow a$ or $\uparrow a$ instead of $\downarrow\{a\}$ or $\uparrow\{a\}$, respectively.
Let $P$ be a poset. We denote $\sigma(P)$, $\upsilon(P)$ and $A(P)$ to be the Scott topology, the upper topology and the Alexandroff topology on $P$ respectively. For a complete lattice $L$, and $x,y\in L$, we say that $x\prec y$ iff $y\in\mathrm{int}_{\upsilon(L)}\uparrow x$. $L$ is hypercontinuous if $\{d\in L:d\prec x\}$ is directed and has $x$ as a supremum for all $x\in L$. $L$ is hyperalgebraic if $x=\bigvee\{d\in L,d\prec d\leq x\}$ for all $x\in L$. An element $y$ such that $y\prec y$ is called hypercompact.
@@
For any two topological spaces $X,Y$, we denote $Y^{X}$ or $TOP(X,Y)$ the set of all continuous maps from $X$ to $Y$, endowed with the pointwise order. Denote $[X\to Y]_{p}$ and $[X\to Y]_{I}$ to be the topological space equipped with the topology of pointwise convergence and the Isbell topology on $Y^{X}$ respectively.
We now introduce the notion of a monotone determined space.
Let $(X,\mathcal{O}(X))$ be a $T_{0}$ space. Every directed subset $D\subseteq X$ can be regarded as a monotone net $(d)_{d\in D}$. Set $DS(X)=\{D\subseteq X:D\ {\rm is\ directed}\}$ to be the family of all directed subsets of $X$. For an $x\in X$, we denote $D\rightarrow x$ to mean that $x$ is a limit of $D$, i.e., $D$ converges to $x$ with respect to the topology on $X$. Then the following result is obvious.
@@
Let $X$ be a $T_{0}$ space. For any $(D,x)\in DS(X)\times X$, $D\rightarrow x$ if and only if $D\cap U\not=\emptyset$ for any open neighborhood of $x$.
Set $DLim(X)=\{(D,x)\in DS(X)\times X:\ D\rightarrow x\}$ to be the set of all pairs of directed subsets and their limits in $X$. Then ($(\{y\},x)\in DLim(X)$ iff $x\sqsubseteq y$ for all $x,y\in X$.
@@
Let $X$ be a $T_{0}$ space. A subset $U\subseteq X$ is called directed-open if for all $(D,x)\in DLim(X)$, $x\in U$ implies $D\cap U\not=\emptyset$.
@@
[24] Let $X$ be a $T_{0}$ topological space. Then
@@
    $X$ equipped with $d(\mathcal{O}(X))$ is a $T_{0}$ topological space such that $\sqsubseteq_{d}=\sqsubseteq$, where $\sqsubseteq_{d}$ is the specialization order relative to $d(\mathcal{O}(X))$.
@@
[24] A topological space $X$ is said to be a monotone determined space if it is $T_{0}$ and every directed-open set is open; equivalently, $d(\mathcal{O}(X))=\mathcal{O}(X)$.
@@
[24] Let $X$ be a $T_{0}$ space, we have
@@
      For all $U\subseteq X$, $U$ is open iff for any $(D,x)\in DLim(X)$, $x\in U$ implies $U\cap D\not=\emptyset$.
@@
      For all $A\subseteq X$, $A$ is closed iff for any directed subset $D\subseteq A$, $D\rightarrow x$ implies $x\in A$ for all $x\in X$.
@@
    ([8]) Every poset endowed with the weak Scott topology is a monotone determined space.
@@
Given any two monotone determined spaces $X,Y$, a subset $U$ of $X\otimes Y$ is open iff the following conditions hold:
@@
    for any directed subsets $(x_{i})_{I}$ with $(x_{i})_{I}\to x$ in $X$ and any $(x,y)\in U$, we have $((x_{i},y))_{I}\cap U\neq\emptyset$;
@@
    for any directed subsets $(y_{i})_{I}$ with $(y_{i})_{I}\to y$ in $Y$ and any $(x,y)\in U$ we have $((x,y_{i}))_{I}\cap U\neq\emptyset$.
Proof. Assume that $U$ is open in $X\otimes Y$ and $(x,y)\in U$. For any directed set $(x_{i})_{I}\to x$ in $X$ and any $y\in Y$, $((x_{i},y))_{I}$ is a directed subset of $X\otimes Y$ and $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$. By Theorem 2.3 (3), $((x_{i},y))_{I}\to(x,y)$ in $X\otimes Y$. Thus, $((x_{i},y))_{I}\cap U\neq\emptyset$. It is the same for $(y_{i})_{I}$.
Conversely, assume that $(1)$ and $(2)$ are satisfied. We show that $U$ is open in $X\otimes Y$. It is easily seen that $U$ is an upper set relative to the specialization order of $X\times Y$. Let $D=((x_{i},y_{i}))_{I}$ be a directed subset of $X\times Y$ and converges to $(x,y)\in U$ in $X\times Y$. We have that $(x_{i})_{I}\to x$ in $X$ and $(y_{i})_{I}\to y$ in $Y$ respectively. Thus, $((x_{i},y))_{I}\to(x,y)$ in $X\times Y$, and then there exists some $i_{0}\in I$ such that $(x_{i_{0}},y)\in U$. By $((x_{i_{0}},y_{i}))_{I}\to(x_{i_{0}},y)$, there exists some $i_{1}\in I$ such that $(x_{i_{0}},y_{i_{1}})\in U$. Let $i_{0},i_{1}\leq i_{2}$, then $(x_{i_{2}},y_{i_{2}})\in U$. $\Box$
3 D-completion of monotone determined spaces
Monotone convergence spaces (a.k.a d-spaces) is a very important class of topological spaces in domain theory. It is well known that every $T_{0}$ space has a D-completion. Let us recall some results about D-completion for further discussions.
Definition 3.8.
Let $X$ be a $T_{0}$ space. If $X$ with the specialization order is a dcpo and every open set of $X$ is Scott open in $(X,\leq)$, then we call $X$ a monotone convergence space space.
A subset $A$ of a poset $P$ is called d-closed if for every directed set $D$ of $A$ that possesses supremum $\bigvee D$, $\bigvee D$ is included in $A$. The d-closed sets form the closed sets for a topology, called the d-topology [16]. We denote the closure of $A$ in $P$ with d-topology by $\text{cl}_{d}(A)$. If $\text{cl}_{d}(A)$ is equal to $P$, then we say that $A$ is d-dense in $P$. A map $f$ from a poset $P$ to a poset $Q$ is called d-continuous if $f$ is continuous under the d-topology.
Definition 3.9.
[16] An embedding $j\colon X\rightarrow\tilde{X}$ of a space $X$ with image a d-dense subset of a d-space $\tilde{X}$ is called a $D$-completion.
Lemma 3.10.
[16] Consider the following properties for a subset $A$ of a $T_{0}$ space
@@ 3 D-completion of monotone determined spaces
    $A$ is a monotone convergence space.
  2. (2)
    $A$ is a sub-dcpo.
  3. (3)
    $A$ is d-closed.
@@
[16] Let $j\colon X\rightarrow Y$ be a topological embedding of a space $X$ into a monotone convergence space $Y$. Let $\tilde{X}=\text{cl}_{d}(j(X))$ be the d-closure of $j(X)$ in $Y$, equipped with the relative topology from $Y$. Then $k\colon X\rightarrow\tilde{X}$, the corestriction of $j$, is a universal $D$-completion, that is, for every continuous map $f$ from $X$ to a monotone convergence space $M$, there is a unique continuous map $\tilde{f}\colon\tilde{X}\rightarrow M$ such that $\tilde{f}\circ k=f$.
@@
For any $T_{0}$ space $X$, there is a natural monotone convergence space $C(X)$ (the lattice of closed subsets of $X$) with the upper topology such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow C(X)$ is an embedding. We call the d-closure of $\eta(X)$ in $C(X)$ the standard D-completion of $X$ and denote it by $X^{d}$.
@@
Let $X$ be a $T_{0}$ space and $X^{d}$ be its D-completion. Then $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$.
@@
[…]
@@
Let $D$ be a set ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$. The order on $D$ is defined as follows: $x\leq y$ iff
@@
$a_{1}$$a_{2}$$a_{3}$$a_{3}$$\top$
Figure 1: ×$\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$

### v2

@@ Abstract
We present several equivalent conditions of the continuity of the supremum function from the square of the Scott space of $C(X)$ to itself under mild assumptions, where $C(X)$ denotes the lattice of closed subsets of a $\mathbf{T_{0}}$ topological space. We also show that a $\mathbf{T_{0}}$ space is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets is a quasicontinuous (quasialgebraic) domain by using $n$-approximation. Furthermore, we provide a necessary condition for when a topological space possesses a Scott completion. This allows us to give more examples which do not have Scott completions.
@@ 1 Introduction
Topological semilattice are closely connected to the theory of continuous lattices in domain theory. For instance, Lawson showed that compact unital semilattices with small semilattices are exactly continuous lattices equipped with the Lawson topology [11, Theorem VI-3.4]. From the viewpoint of category topology, Banaschewski introduced the notion of an essential embedding within the category of $\mathbf{T_{0}}$ spaces and continuous maps [2], which has a close relation to Scott’s injective spaces. He also obtained a characterization of so-called essentially complete $\mathbf{T_{0}}$ spaces in terms of join filters. Subsequently, Hoffmann provided a more transparent characterization: essentially complete $\mathbf{T_{0}}$ spaces are precisely the sober spaces that form complete lattices under their specialization order, satisfying an additional condition relating the supremum operator to the topology [14]. Powerdomains play an important role in modeling the semantics of nondeterministic programming languages (see [1, Chapter 6] and [11, IV-8]). The Hoare powerdomain over a topological space can be realized as the set of its nonempty closed subsets ordered by inclusion. In this paper we study a special kind of essentially complete $\mathbf{T_{0}}$ space, namely the Scott space of the lattice of closed subsets of a topological space with the supremum operator.
Directed spaces [24] constitute a special class of $\mathbf{T_{0}}$ topological spaces which extend the concept of Scott spaces. Within $\mathbf{T_{0}}$ spaces, the notion of directed spaces is equivalent both to monotone determined spaces (as established by Erné [8]) and to CRP-spaces (introduced by Battenfeld [3, Definition 6.7]). $\mathbf{T_{0}}$ c-spaces (or $n$-continuous spaces) are all monotone determined spaces. Based on this, Feng and Kou [9] introduced another approximation relation on monotone determined spaces, called $d$-approximation, and defined the quasicontinuous monotone determined spaces. They established an equivalence between quasicontinuous spaces and locally hypercompact spaces. In domain theory, there is a close relationship between the continuity of a dcpo $L$ and that of the lattice $\sigma(L)$ of its Scott-open subsets. More surprisingly, the continuity of a dcpo $L$ can be characterized by the continuity of the lattice $\Gamma(L)$ of its Scott-closed subsets (see [4, Theorem 3.17]). Moreover, this correspondence extends to general $\mathbf{T_{0}}$ spaces. We show that a $\mathbf{T_{0}}$ space $X$ is quasicontinuous (quasialgebraic) iff the lattice of its closed subsets $C(X)$ is quasicontinuous (quasialgebraic) by using $n$-approximation.
@@
It is well known that for completely distributive (resp., distributive hypercontinuous) lattices $L$, the hull-kernel topology on the spectrum $\mathrm{Spec}L$ coincides with the Scott topology [11]. A longstanding open problem posed by Lawson and Mislove asks: for which distributive continuous lattices does the spectrum, equipped with the hull-kernel topology, form precisely a sober locally compact Scott space? (See [23, Problem 528].) Chen, Kou, and Lyu [4] established a necessary condition for this problem: $\sigma(L^{\mathrm{op}})=v(L^{\mathrm{op}})$. Since $L^{\mathrm{op}}$ corresponds to the lattice of closed subsets of a topological space, we investigate equivalent conditions for $\sigma(C(X))=v(C(X))$ in general topological spaces $X$. Surprisingly, the continuity of the supremum map $\sup\colon\Sigma C(X)\times\Sigma C(X)\to\Sigma C(X)$ constitutes one such equivalent condition under mild assumptions. For monotone determined spaces, the additional assumption can be dropped. In addition, we provide an example of non-monotone determined space $X$ such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow\Sigma C(X)$ is continuous.
@@
Monotone convergence spaces play a fundamental role in domain theory. Every $\mathbf{T_{0}}$ space admits a free monotone convergence completion, known as its D-completion [16]. We show that D-completion of any monotone determined space is a Scott space. The concept of Scott completion, introduced by Zhang, Shi, and Li [25], refers to a construction that slightly differs from D-completion. As a corollary, we establish that for any monotone determined space, its Scott completion coincides with its D-completion. The Scott completion of monotone determined spaces are shown to be very useful to construct free dcpo algebras over any general dcpo (see [6, 7]). But the notion does not seem to have been studied much. We provide a necessary condition for when a space has Scott completion. Based on this we give a simple proof of a previous result in [17, Theorem 4.1].
@@
Outline. In Section 2, we recall some useful knowledge about monotone determined spaces. We discuss D-completion of monotone determined spaces and provide more examples of topological spaces which do not have Scott completions are presented in Section 3. In Section 4 the Scott space of the lattice of closed subsets of a topological space with the jointly continuous supremum operator is discussed. We characterize the quasicontinuity of lattice of closed subsets in Section 5.
@@ 2 Preliminaries
A nonempty set $P$ endowed with a partial order $\leq$ is called a poset. For $A\subseteq P$, we set $\downarrow\hskip-2.0ptA=\{x\in P:\exists a\in A,\ x\leq a\}$$\uparrow\hskip-2.0ptA=\{x\in P:\exists a\in P,\ a\leq x\}$AA is called a lower or an upper set, if $A=\downarrow\hskip-2.0ptA$ or $A=\uparrow\hskip-2.0ptA$ respectively. For an element $a\in P$, we use $\downarrow\hskip-2.0pta$ or $\uparrow\hskip-2.0pta$ instead of $\downarrow\hskip-2.0pt\{a\}$ or $\uparrow\hskip-2.0pt\{a\}$, respectively.
Let $P$ be a poset. We denote $\sigma(P)$, $\upsilon(P)$ and $A(P)$ to be the Scott topology, the upper topology and the Alexandroff topology on $P$ respectively. For a complete lattice $L$, and $x,y\in L$, we say that $x\prec y$ iff $y\in\mathrm{int}_{\upsilon(L)}\uparrow\hskip-2.0ptx$. $L$ is hypercontinuous if $\{d\in L:d\prec x\}$ is directed and has $x$ as a supremum for all $x\in L$. $L$ is hyperalgebraic if $x=\bigvee\{d\in L:d\prec d\leq x\}$ for all $x\in L$. An element $y$ such that $y\prec y$ is called hypercompact.
@@
For any two topological spaces $X,Y$, we denote $Y^{X}$ or $TOP(X,Y)$ the set of all continuous maps from $X$ to $Y$, endowed with the pointwise order. Denote $X\to Y]_{p}$ and $X\to Y]_{I}$ to be the topological space equipped with the topology of pointwise convergence and the Isbell topology on $Y^{X}$ respectively.
We now introduce the notion of a monotone determined space. Let $(X,\mathcal{O}(X))$ be a $\mathbf{T_{0}}$ space. Every directed subset $D\subseteq X$ can be regarded as a monotone net $(d)_{d\in D}$. Set $DS(X)=\{D\subseteq X:D\ {\rm is\ directed}\}$ to be the family of all directed subsets of $X$. For an $x\in X$, we denote $D\rightarrow x$ to mean that $x$ is a limit of $D$, i.e., $D$ converges to $x$ with respect to the topology on $X$. Then the following result is obvious.
@@
Let $X$ be a $\mathbf{T_{0}}$ space. For any $(D,x)\in DS(X)\times X$, $D\rightarrow x$ if and only if $D\cap U\not$$\emptyset$ for any open neighborhood of $x$.
Set $DLim(X)=\{(D,x)\in DS(X)\times X:\ D\rightarrow x\}$ to be the set of all pairs of directed subsets and their limits in $X$. Then $(\{y\},x)\in DLim(X)$ iff $x\sqsubseteq y$ for all $x,y\in X$.
@@
Let $X$ be a $\mathbf{T_{0}}$ space. A subset $U\subseteq X$ is called directed-open if for all $(D,x)\in DLim(X)$, $x\in U$ implies $D\cap U\not$$\emptyset$.
@@
[24, Theorem 2.4] Let $X$ be a $\mathbf{T_{0}}$ topological space. Then
@@
    $X$ equipped with $d(\mathcal{O}(X))$ is a $\mathbf{T_{0}}$ topological space such that $\sqsubseteq_{d}=\sqsubseteq$, where $\sqsubseteq_{d}$ is the specialization order relative to $d(\mathcal{O}(X))$.
@@
[24, Definition 3.1] A topological space $X$ is said to be a monotone determined space if it is $\mathbf{T_{0}}$ and every directed-open set is open; equivalently, $d(\mathcal{O}(X))=\mathcal{O}(X)$.
@@
[24, Theorem 3.2] Let $X$ be a $\mathbf{T_{0}}$ space, we have
@@
      For all $\subseteq X, UU$ is open iff for any $(D,x)\in DLim(X)$, $x\in U$ implies $U\cap D\not$$\emptyset$.
@@
      For all $\subseteq X, AA$ is closed iff for any directed subset $D\subseteq A$, $D\rightarrow x$ implies $x\in A$ for all $x\in X$.
@@
    ([8, Section 7]) Every poset endowed with the weak Scott topology is a monotone determined space.
@@
3 D-completion and Scott completion
Monotone convergence spaces (a.k.a d-spaces) is a very important class of topological spaces in domain theory. It is well known that every $\mathbf{T_{0}}$ space has a D-completion. Let us recall some results about D-completion for further discussions.
Definition 3.7.
Let $X$ be a $\mathbf{T_{0}}$ space. If $X$ with the specialization order is a dcpo and every open set of $X$ is Scott open in $(X,\leq)$, then we call $X$ a monotone convergence space space.
A subset $A$ of a poset $P$ is called d-closed if for every directed set $D$ of $A$ that possesses supremum $\bigvee D$, $\bigvee D$ is included in $A$. The d-closed sets form the closed sets for a topology, called the d-topology [16, Section 5]. We denote the closure of $A$ in $P$ with d-topology by $\text{cl}_{d}(A)$. If $\text{cl}_{d}(A)$ is equal to $P$, then we say that $A$ is d-dense in $P$. A map $f$ from a poset $P$ to a poset $Q$ is called d-continuous if $f$ is continuous under the d-topology.
Definition 3.8.
[16, Definition 6.5] An embedding $j\colon X\rightarrow\tilde{X}$ of a space $X$ with image a d-dense subset of a d-space $\tilde{X}$ is called a $D$-completion.
@@ 3 D-completion and Scott completion
[16, Lemma 6.3] Consider the following properties for a subset $A$ of a $\mathbf{T_{0}}$ space
@@
    AA is a monotone convergence space.
@@
    AA is a sub-dcpo.
@@
    AA is d-closed.
@@
[16, Theorem 6.7] Let $j\colon X\rightarrow Y$ be a topological embedding of a space $X$ into a monotone convergence space $Y$. Let $\tilde{X}=\text{cl}_{d}(j(X))$ be the d-closure of $j(X)$ in $Y$, equipped with the relative topology from $Y$. Then $k\colon X\rightarrow\tilde{X}$, the corestriction of $j$, is a universal $D$-completion, that is, for every continuous map $f$ from $X$ to a monotone convergence space $M$, there is a unique continuous map $\tilde{f}\colon\tilde{X}\rightarrow M$ such that $\tilde{f}\circ k=f$.
@@
For any $\mathbf{T_{0}}$ space $X$, there is a natural monotone convergence space $C(X)$ (the lattice of closed subsets of $X$) with the upper topology such that $\eta=\lambda x.{\downarrow}x\colon X\rightarrow C(X)$ is an embedding. We call the d-closure of $\eta(X)$ in $C(X)$ the standard D-completion of $X$ and denote it by $X^{d}$.
@@
Let $X$ be a $\mathbf{T_{0}}$ space and $X^{d}$ be its D-completion. Then $\mathcal{O}(X)\cong\mathcal{O}(X^{d})$.
@@
$A\in X^{d}=\text{cl}_{d}\{{\downarrow}x:x\in X\}=\text{cl}_{d}\{{\downarrow}x:x\in A\}\cup\text{cl}_{d}\{{\downarrow}x:x\in X\backslash A\}.$
@@
Let $D$ be a set $\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$. The order on $D$ is defined as follows: $x\leq y$ iff
@@
$D$ can be easily depicted as in Figure 1. Let $X=\mathbb{N}\times\mathbb{N}\cup\{\top\}$ with the subspace topology of $\Sigma D$. It is easy to see that $\Sigma D$ is a D-completion of $X$. Notice that $\{\top\}$ is always a directed open set of $X$ which is not open in $X$. Hence $X$ is not a monotone determined space.
$a_{1}$$a_{2}$$a_{3}$$a_{3}\top$
Figure 1: $\mathbb{N}\times\mathbb{N}\cup\{a_{n}:n\in\mathbb{N}\}\cup\{\top\}$
The concept of Scott completion is introduced by Zhang, Shi and Li in [25]. It has played an important role in constructing free dcpo algebras over general dcpos, see [4, 6]. But the notion does not seem to have been investigated much. We will give more examples of topological spaces which do not have Scott completions in this section.
Definition 3.17.
[25] A Scott completion $(Y,f)$ of a space $X$ is a Scott space $Y$ together with a continuous map $f\colon X\rightarrow Y$ such that for any Scott space $Z$ and continuous map $g\colon X\rightarrow Z$, there exists a unique continuous map $\tilde{g}$ satisfying $g=\tilde{g}\circ f$.
Theorem 3.18.
@@
Proof.
Let $Y\rightarrow 2]$ denote the set of continuous functions from $Y$ to the Sierpinski space $2$ with the pointwise order. It is well known that $\mathcal{O}(Y)\cong[Y\rightarrow 2]$. For any $f\in[X\rightarrow 2]$, there exists only one $\bar{f}\in[\tilde{X}\rightarrow 2]$ such that $f=\tilde{f}\circ i$ since $(\tilde{X},i)$ is a Scott-completion of $X$. Hence we obtain a map $\theta\colon[X\rightarrow 2]\rightarrow[\tilde{X}\rightarrow 2]$ as follows
$\theta(f)=\bar{f}.$
Clearly $\theta$ is bijective.
It is easy to see that $\theta$ reflects order. Hence $\theta$ is an order isomorphism. So we obtain that $𝒪(X~)\mathcal$$\cong\mathcal{O}(\tilde{X})$. ∎
Remark 3.19.
Corollary 3.20.
  1. (1)
  2. (2)
    Every sober but non-Scott space does not have a Scott-completion.
Proof.
[…]
The following result is a direct application of Theorem 3.18.
Corollary 3.21.
Proof.
[…]
Proposition 3.22.
Let $X$ be a $\mathbf{T_{0}}$ space such that it has a Scott-completion $(\tilde{X},i)$. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
Proof.
[…]
$\eta=i^{-1}\circ j\circ i.$
Corollary 3.23.
Let $X$ be a $\mathbf{T_{0}}$ topological space such that its D-completion is a Scott space. Then $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.
Example 3.24.
There exists a topological space $X$ such that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous while $X$ does not have a Scott completion. Let $\mathcal{J}$ be Johnstone’s non-sober dcpo [12, Exercise 5.2.15]. Notice that $Id(\prod^{n}\mathcal{J})=\prod^{n}Id(\mathcal{J})$}, by using Corollary 4.32 and Proposition 4.35 we obtain that $\sigma(\Gamma(\mathcal{J}))=v(\Gamma(\mathcal{J}))$. Now let $X$ be the soberification of $\Sigma\mathcal{J}$. It implies that $\sigma(C(X))=v(C(X))$. Since $X$ is not a Scott space, $X$ does not have a Scott completion by Corollary 3.20. It is easy to verify that $\eta\colon X\rightarrow\Sigma C(X)$ is continuous.